# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ B.1 Proof of Theorem 1
$\widehat{Cov}(x_{i},x_{j})\stackrel{{\scriptstyle d}}{{\rightarrow}}N\mathopen{}\mathclose{{}\left(0\>,\>\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)\ \ .$
Consider that $E[a_{ii}]=\frac{T-1}{1-\phi_{i}^{2}}$ (see Lemma B.2). Then, $b=\frac{a_{ji}}{a_{ii}}$ is Normally distributed and, based on the approximation of mean and variance of a ratio (see  Stuart and Ord, 1998), we have $E[b]=0$ and
$Var[b]=(T-1)^{2}\frac{Var[a_{ij}]}{E[a_{ii}]^{2}}=\frac{(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{i}^{2})}{(T-1)(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}\ \ .$
$\blacksquare$
Lemma B.1 shows that the OLS estimate $b$ is normally distributed with a variance that strongly depends on the degree of predictors serial correlation. In the context of finite sample, it is common to adjust the standard error of the OLS to achieve consistency in the presence of heteroskedasticity and/or serial correlation; this leads, for instance, to the Heteroskedasticity and Autocorrelation Consistent (HAC) estimator of Newey and West (1987) (NW). However, NW estimates can be highly sub-optimal (or inefficient) in the presence of strong serial correlation (Baillie et al., 2024). In Supplement J we provide a simulation study to validate the approximation of the sample distribution of $b$ to the probability distribution in Lemma B.1.
Probability Distribution of $v$.
Here we derive the probability distribution of the sum of the square of residuals obtained by regressing $x_{j}$ on $x_{i}$. Since $v=a_{jj}-a_{ji}^{2}/a_{ii}$, we start by deriving the distribution of $a_{jj}$ and $a_{ji}^{2}/a_{ii}$ in the following two Lemmas.
Lemma B.2:
Let $\xi_{a}=\mathopen{}\mathclose{{}\left[3(T-1)-(T-1)^{2}+2\sum_{t=1}^{T-2}(1+2\phi_{j}^{2t})}\right]$. Then, $a_{jj}\stackrel{{\scriptstyle d}}{{\rightarrow}}\Gamma\mathopen{}\mathclose{{}\left(\frac{(T-1)^{2}}{\xi_{a}},\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}}\right)$.
Proof: Let $z_{j,t}$ be the variable obtained by standardizing $x_{j,t}$, and observe that $a_{jj}=\sum_{t=1}^{T-1}\frac{z_{j,t}^{2}}{1-\phi_{j}^{2}}$. Thus, $a_{jj}$ is the sum of $T-1$ correlated $\chi_{1}^{2}$ multiplied by $\frac{1}{1-\phi_{j}^{2}}$, which is the case of a Gamma distribution with shape parameter $k_{a}$ and a scale parameter $\theta_{a}$. We have $E(a_{jj})=\frac{T-1}{1-\phi_{j}^{2}}$ and, consequently to the dependency between the elements of $a_{jj}$, $Var(a_{jj})=\xi_{a}(1-\phi_{j}^{2})^{-2}$, where $\xi_{a}=\mathopen{}\mathclose{{}\left[3(T-1)-(T-1)^{2}+2\sum_{t=1}^{T-2}(1+2\phi_{j}^{2t})}\right]$. We can use these moments to obtain $k_{a}=\frac{E(a_{jj})^{2}}{Var(a_{jj})}=\frac{(T-1)^{2}}{\xi_{a}}$ and $\theta_{a}=\frac{Var(a_{jj})}{E(a_{jj})}=\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}$. Therefore $a_{jj}\stackrel{{\scriptstyle d}}{{\rightarrow}}\Gamma\mathopen{}\mathclose{{}\left(\frac{(T-1)^{2}}{\xi_{a}},\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}}\right)$. $\blacksquare$
Lemma B.3:
$a_{ij}^{2}/a_{ii}\stackrel{{\scriptstyle d}}{{\rightarrow}}\Gamma\mathopen{}\mathclose{{}\left(\frac{1}{2},\frac{2(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)$.
Proof: Note that $a_{ij}/\sqrt{a_{ii}}=\sqrt{a_{ii}}b$. Thus, by Lemma B.1 we have that $a_{ij}/\sqrt{a_{ii}}=\sqrt{a_{ii}}b\stackrel{{\scriptstyle d}}{{\rightarrow}}N\mathopen{}\mathclose{{}\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)$. Let $z$ be the variable obtained by standardizing $a_{ij}/\sqrt{a_{ii}}$, we have $a_{ij}^{2}/a_{ii}=\frac{z^{2}(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ where $E(a_{ij}^{2}/a_{ii})=\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ and $Var(a_{ij}^{2}/a_{ii})=2\mathopen{}\mathclose{{}\left(\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)^{2}$. Using the same argument as in Lemma B.2, we obtain $a_{ij}^{2}/a_{ii}\stackrel{{\scriptstyle d}}{{\rightarrow}}\Gamma\mathopen{}\mathclose{{}\left(\frac{1}{2},\frac{2(1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)$. $\blacksquare$
Lemmas B.2 and B.3 allow us to derive the finite sample distribution of $v$.
Lemma B.4:
Let $T_{v}=\mathopen{}\mathclose{{}\left\lceil\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}\right\rfloor$, and $\xi_{v}=\mathopen{}\mathclose{{}\left[3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(1+2\phi_{j}^{2t})}\right]$. Then, $v=a_{jj}-a_{ji}^{2}/a_{ii}\stackrel{{\scriptstyle d}}{{\rightarrow}}\Gamma\mathopen{}\mathclose{{}\left(\frac{T_{v}^{2}}{\xi_{v}},\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}}\right)$.
Proof: We combine the results in Lemmas B.2 and B.3. Considering $E(v)=E(a_{jj}-a_{ij}^{2}/a_{ii})=E(a_{jj})-E(a_{ij}^{2}/a_{ii})=\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{1}^{2}\phi_{2}^{2})}{(1-\phi_{2}^{2})(1-\phi_{i}\phi_{j})^{2}}$, we define $T_{v}=\mathopen{}\mathclose{{}\left\lceil\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}\right\rfloor$. Therefore, by using the same argument as in Lemma B.2, we have that $v$ approximates the sum of $T_{v}$ scaled $\chi^{2}_{1}$. A brief algebraic manipulation reveals that the variance of $v$ is $\xi_{v}(1-\phi_{jj})^{-2}$, where $\xi_{v}=\mathopen{}\mathclose{{}\left[3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(1+2\phi_{j}^{2t})}\right]$ incorporates the dependence between $x_{j,t}$’s. Therefore, we have that $v=a_{jj}-a_{ji}^{2}/a_{ii}\stackrel{{\scriptstyle d}}{{\rightarrow}}\Gamma\mathopen{}\mathclose{{}\left(\frac{T_{v}^{2}}{\xi_{v}},\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}}\right)$. $\blacksquare$
Probability Density of $\widehat{c}_{ij}^{x}$.
Note that $\sqrt{a_{ii}}b$ and $v$ are independent. Using Lemmas B.1 and B.4 and equation (B.7) we can now derive the probability density of $\widehat{c}_{ij}^{x}$.
Because of Lemma B.1, $\sqrt{a_{ii}}b\stackrel{{\scriptstyle d}}{{\rightarrow}}N\mathopen{}\mathclose{{}\left(0,\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)$. Let $\delta^{2}=\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$, $k_{v}=\frac{T_{v}^{2}}{\xi_{v}}$, $\theta_{v}=\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}$ and $t=\frac{\sqrt{a_{ii}}b}{\sqrt{v}}$. Thus, we have the densities
@@
$g\mathopen{}\mathclose{{}\left(\sqrt{a_{ii}}b}\right)$ $=$ $\frac{1}{\delta\sqrt{2\pi}}\text{exp}\mathopen{}\mathclose{{}\left(-\frac{a_{ii}b^{2}}{2\delta^{2}}}\right)\ ,$
@@
$h(v)$ $=$ $\frac{1}{(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{}\left(k_{v}}\right)}v^{k_{v}-1}\text{exp}\mathopen{}\mathclose{{}\left(-\frac{v}{\theta_{v}}}\right)\ .$
@@
[…]
@@
[…]
The integral on the right-hand side can be represented by using the gamma function
[…]
Thus we obtain
$f(t)$ $=$ $\Upsilon\ \mathopen{}\mathclose{{}\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}\right)^{-(k_{v}+\frac{1}{2})}\Gamma\mathopen{}\mathclose{{}\left(k_{v}+\frac{1}{2}}\right)$
$=$ $\Upsilon\ \mathopen{}\mathclose{{}\left(\frac{2\delta^{2}+\theta_{v}t^{2}}{\theta_{v}2\delta^{2}}}\right)^{-(k_{v}+\frac{1}{2})}\Gamma\mathopen{}\mathclose{{}\left(k_{v}+\frac{1}{2}}\right)$
$=$ $\frac{\Gamma\mathopen{}\mathclose{{}\left(k_{v}+\frac{1}{2}}\right)}{\sqrt{2\pi}\delta(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{}\left(k_{v}}\right)}\mathopen{}\mathclose{{}\left(\frac{2\delta^{2}+\theta_{v}t^{2}}{\theta_{v}2\delta^{2}}}\right)^{-(k_{v}+\frac{1}{2})}$
$=$ $\frac{\Gamma\mathopen{}\mathclose{{}\left(k_{v}+\frac{1}{2}}\right)\sqrt{\theta_{v}}}{\delta\sqrt{2\pi}\Gamma\mathopen{}\mathclose{{}\left(k_{v}}\right)}\mathopen{}\mathclose{{}\left[\frac{2\delta^{2}+\theta_{v}t^{2}}{2\delta^{2}}}\right]^{-(k_{v}+\frac{1}{2})}.$
@@
$=$ $\frac{\Gamma\mathopen{}\mathclose{{}\left(k_{v}+\frac{1}{2}}\right)(1-\phi_{i}\phi_{j})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{}\left(k_{v}}\right)\sqrt{2\pi T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}\mathopen{}\mathclose{{}\left(1+\frac{t^{2}\xi_{v}(1-\phi_{i}\phi_{j})^{2})}{2T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}\right)^{-(k_{v}+\frac{1}{2})}\ .$
@@
$=$ $\Theta\mathopen{}\mathclose{{}\left[1+\frac{r^{2}(1-\ddot{\phi})^{2}\xi_{v}}{(1-r^{2})2T_{v}(1-\ddot{\phi}^{2})}}\right]^{-(k_{v}+\frac{1}{2})}\mathopen{}\mathclose{{}\left[1-r^{2}}\right]^{-\frac{3}{2}}$
$=$ $\Theta\mathopen{}\mathclose{{}\left[\frac{(1-r^{2})2T_{v}(1-\ddot{\phi}^{2})+r^{2}\xi_{v}(1-\ddot{\phi})^{2}}{(1-r^{2})2T_{v}(1-\ddot{\phi}^{2})}}\right]^{-(k_{v}+\frac{1}{2})}\mathopen{}\mathclose{{}\left[1-r^{2}}\right]^{-\frac{3}{2}}$
@@
Thus, the probability density of $\widehat{c}_{ij}^{x}$ is
@@
$\blacksquare$
B.2 Proof of Theorem 2

### v3

@@ A.1 Proof of Proposition 1
$\widehat{Cov}(x_{i,t},x_{j,t})\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0\>,\>\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)\ \ .$
To obtain a closed–form representation of the distribution of $b=a_{ji}/a_{ii}$, we adopt a plug–in approximation by replacing $a_{ii}$ with its expectation $E[a_{ii}]=(T-1)/(1-\phi_{i}^{2})$. Following standard approximations for ratios of random variables (see Stuart1998), we have
@@
$b\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{i}^{2})}{(T-1)(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)\ \ .$
$\square$
Distribution of $v$. Here, we derive the sample distribution of the sum of the square of residuals obtained by regressing $x_{j}$ on $x_{i}$. Since $v=a_{jj}-a_{ji}^{2}/a_{ii}$, we start by deriving the distribution of $a_{jj}$ and $a_{ji}^{2}/a_{ii}$ in the following two Lemmas.
Lemma 2:
@@
The sample probability distribution of $a_{jj}$ is approximately $\Gamma\mathopen{}\mathclose{{\left(\frac{(T-1)^{2}}{\xi_{a}},\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}}}\right)$, where the quantity $\xi_{a}=\mathopen{}\mathclose{{\left[3(T-1)-(T-1)^{2}+2\sum_{t=1}^{T-2}(T-1-t)(1+2\phi_{j}^{2t})}}\right]$.
Proof of Lemma 2 Let $z_{j,t}$ be the standardized version of $x_{j,t}$, so that $x_{j,t}=z_{j,t}/\sqrt{1-\phi_{j}^{2}}$. Then $a_{jj}=\sum_{t=1}^{T}(x_{j,t}-\bar{x}_{j})^{2}\;\approx\;\sum_{t=1}^{T}x_{j,t}^{2}=\frac{1}{1-\phi_{j}^{2}}\sum_{t=1}^{T}z_{j,t}^{2}$. Using the T1T-1 normalization for sample variances and covariances, we approximate the above quadratic form by a sum of T1T-1 correlated $\chi^{2}_{1}$ terms, that is, we work with an effective number of degrees of freedom equal to T1T-1. Thus, $a_{jj}$ is approximated by a Gamma distribution with shape parameter $k_{a}$ and scale parameter $\theta_{a}$, obtained by matching the first two moments. Thus, $a_{jj}$ is the sum of T1T-1 correlated $\chi_{1}^{2}$ multiplied by $\frac{1}{1-\phi_{j}^{2}}$, approximate a Gamma distribution with shape parameter $k_{a}$ and a scale parameter $\theta_{a}$. Thus, we have to define such parameters via moments matching. We have $E(a_{jj})=\frac{T-1}{1-\phi_{j}^{2}}$ and, consequently to the dependency between the elements of $a_{jj}$, $Var(a_{jj})=\xi_{a}(1-\phi_{j}^{2})^{-2}$, where $\xi_{a}=\mathopen{}\mathclose{{\left[3(T-1)-(T-1)^{2}+2\sum_{t=1}^{T-2}(T-1-t)(1+2\phi_{j}^{2t})}}\right]$. We can use these moments to obtain $k_{a}=\frac{E(a_{jj})^{2}}{Var(a_{jj})}=\frac{(T-1)^{2}}{\xi_{a}}$ and $\theta_{a}=\frac{Var(a_{jj})}{E(a_{jj})}=\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}$. Therefore $a_{jj}\approx\Gamma\mathopen{}\mathclose{{\left(\frac{(T-1)^{2}}{\xi_{a}},\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}}}\right)$. $\square$
Lemma 3:
The sample probability distribution of $a_{ij}^{2}/a_{ii}$ is approximately $\Gamma\mathopen{}\mathclose{{\left(\frac{1}{2},\frac{2(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$.
Proof of Lemma 3 Note that $a_{ij}/\sqrt{a_{ii}}=\sqrt{a_{ii}}b$. Thus, by Lemma 1 we have that $a_{ij}/\sqrt{a_{ii}}=\sqrt{a_{ii}}b\approx N\mathopen{}\mathclose{{\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$. Let $z$ be the variable obtained by standardizing $a_{ij}/\sqrt{a_{ii}}$, we have $a_{ij}^{2}/a_{ii}=\frac{z^{2}(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ where $E(a_{ij}^{2}/a_{ii})=\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ and $Var(a_{ij}^{2}/a_{ii})=2\mathopen{}\mathclose{{\left(\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)^{2}$. Using the same argument as in Lemma 2, we obtain $a_{ij}^{2}/a_{ii}\approx\Gamma\mathopen{}\mathclose{{\left(\frac{1}{2},\frac{2(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$. $\square$
Lemmas 2 and 3 allow us to derive the sample distribution of $v$.
Lemma 4:
The sample probability distribution of $v=a_{jj}-a_{ji}^{2}/a_{ii}$ is approximately $\Gamma\mathopen{}\mathclose{{\left(\frac{T_{v}^{2}}{\xi_{v}},\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}}}\right)$, where the quantities $T_{v}=\mathopen{}\mathclose{{\left\lfloor\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}}\right\rceil$, and $\xi_{v}=\mathopen{}\mathclose{{\left[3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(T_{v}-t)(1+2\phi_{j}^{2t})}}\right]$.
Proof of Lemma 4 While Lemmas 23 provide marginal approximations for $a_{jj}$ and $a_{ji}^{2}/a_{ii}$, the exact finite–sample law of $v=a_{jj}-a_{ji}^{2}/a_{ii}$ does not admit a tractable closed-form expression, since it depends on the joint distribution of these two terms and, in particular, on ${Cov}\!\big(a_{jj},a_{ji}^{2}/a_{ii}\big)$. We therefore introduce an additional moment–matching approximation: we model $v$ as a scaled quadratic form with the same dependence structure as $x_{j}$, but with an effective number of degrees of freedom $T_{v}$ chosen to match $E(v)$. The corresponding variance is approximated by that of a sum of $T_{v}$ correlated $\chi^{2}_{1}$ components (with AR(1) dependence parameter $\phi_{j}$), yielding the Gamma law in Lemma 4.
We combine the results in Lemmas 2 and 3. Considering $E(v)=E\!\mathopen{}\mathclose{{\left(a_{jj}-\frac{a_{ij}^{2}}{a_{ii}}}}\right)=E(a_{jj})-E\!\mathopen{}\mathclose{{\left(\frac{a_{ij}^{2}}{a_{ii}}}}\right)=\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}},$ we define $T_{v}=\mathopen{}\mathclose{{\left\lfloor\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}}\right\rceil$. Therefore, using the same moment–matching argument as in Lemma 2, we approximate $v$ by a sum of $T_{v}$ correlated $\chi^{2}_{1}$ terms with the same AR(1) dependence structure as $x_{j}$. Under this effective degrees–of–freedom approximation, we approximate $\mathrm{Var}(v)$ by the variance of a sum of $T_{v}$ correlated $\chi^{2}_{1}$ terms with AR(1) dependence parameter $\phi_{j}$, yielding $\mathrm{Var}(v)\approx\frac{\xi_{v}}{(1-\phi_{j}^{2})^{2}},$ where $\xi_{v}=\mathopen{}\mathclose{{\left[3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(T_{v}-t)(1+2\phi_{j}^{2t})}}\right]$ incorporates the serial dependence among the $\{x_{j,t}\}$. Matching the first two moments then leads to the Gamma approximation $v=a_{jj}-\frac{a_{ji}^{2}}{a_{ii}}\approx\Gamma\!\mathopen{}\mathclose{{\left(\frac{T_{v}^{2}}{\xi_{v}},\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}}}\right).$ $\square$
Note that the distribution of $b$ and $v$ in the case of independent observations (i.e., without serial correlation) are known (see Anderson03, ch, 4). Here, Lemmas 1 and 4 derive the sample distributions of $b$ and $v$ in the case of serial correlation, namely, by taking into account the dependence of the vectors $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$.
Proof of Proposition 1 Although $\mathbf{x}_{j}$ is Gaussian, its temporal dependence implies that, under OLS, $\sqrt{a_{ii}}\,b$ and $v$ are not exactly independent in finite samples. However, in line with our finite–sample approximation framework, we treat their dependence as a second–order effect and approximate them as independent. Using Lemmas 1 and 4 and equation (1) we can now derive the probability density of $\widehat{c}_{ij}^{x}$. Because of Lemma 1, $\sqrt{a_{ii}}$ is approximately $N\mathopen{}\mathclose{{\left(0,\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$. Let $\delta^{2}=\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$, $k_{v}=\frac{T_{v}^{2}}{\xi_{v}}$, $\theta_{v}=\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}$ and $t=\frac{\sqrt{a_{ii}}b}{\sqrt{v}}$. In the remainder of the proof, we consider the distributions of $\sqrt{a_{ii}}$ and $v$ in Lemmas 1 and 4 as exact. Thus, we have the densities
$g\mathopen{}\mathclose{{\left(\sqrt{a_{ii}}b}}\right)$ $=$ $\frac{1}{\delta\sqrt{2\pi}}\text{exp}\mathopen{}\mathclose{{\left(-\frac{a_{ii}b^{2}}{2\delta^{2}}}}\right)\ ,\ \,\ \ h(v)=\frac{1}{(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}v^{k_{v}-1}\text{exp}\mathopen{}\mathclose{{\left(-\frac{v}{\theta_{v}}}}\right)\ .$
@@
$f(t)$ $=$ $\int\sqrt{v}g\mathopen{}\mathclose{{\left(\sqrt{v}t}}\right)h(v)dv=\int_{0}^{\infty}\sqrt{v}\frac{1}{\delta\sqrt{2\pi}}\text{exp}\mathopen{}\mathclose{{\left(-\frac{vt^{2}}{2\delta^{2}}}}\right)\frac{v^{k_{v}-1}\text{exp}\mathopen{}\mathclose{{\left(-\frac{v}{\theta_{v}}}}\right)}{(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}dv$
$=$ $\frac{1}{\sqrt{2\pi}\delta(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}\int_{0}^{\infty}v^{k_{v}-\frac{1}{2}}\text{exp}\mathopen{}\mathclose{{\left(-\mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)v}}\right)dv\ .$
@@
$f(t)$ $=$ $\Upsilon\int_{0}^{\infty}\mathopen{}\mathclose{{\left[x\mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)^{-1}}}\right]^{k_{v}-\frac{1}{2}}\text{exp}\mathopen{}\mathclose{{\left(-x}}\right)dx$
$=$ $\Upsilon\ \mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}\int_{0}^{\infty}x^{k_{v}+\frac{1}{2}-1}\text{exp}\mathopen{}\mathclose{{\left(-x}}\right)dx\ .$
The integral on the right-hand side can be represented by using the gamma function $\Gamma(\alpha)=\int_{0}^{\infty}x^{\alpha-1}\text{exp}\mathopen{}\mathclose{{\left(-x}}\right)dx$. Thus we obtain
$f(t)$ $=$ $\Upsilon\ \mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)=\Upsilon\ \mathopen{}\mathclose{{\left(\frac{2\delta^{2}+\theta_{v}t^{2}}{\theta_{v}2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)$
$=$ $\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)}{\sqrt{2\pi}\delta(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}\mathopen{}\mathclose{{\left(\frac{2\delta^{2}+\theta_{v}t^{2}}{\theta_{v}2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}=\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)\sqrt{\theta_{v}}}{\delta\sqrt{2\pi}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}\mathopen{}\mathclose{{\left[\frac{2\delta^{2}+\theta_{v}t^{2}}{2\delta^{2}}}}\right]^{-(k_{v}+\frac{1}{2})}.$
@@
$=$ $\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)(1-\phi_{i}\phi_{j})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{2\pi T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}\mathopen{}\mathclose{{\left(1+\frac{t^{2}\xi_{v}(1-\phi_{i}\phi_{j})^{2}}{2T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}}\right)^{-(k_{v}+\frac{1}{2})}\ .$
@@
Thus, the (finite) sample probability density of $\widehat{c}_{ij}^{x}$, tacking the densities in Lemmas 1 and 4 as exact, is
@@
$\square$
A.2 Proof of Proposition 2
By lemma 1 and Hastie2015, ch. 11 , for any $i\in\{1,\dots,n\}$ we have that for $T$ sufficiently large $\mathbf{x}_{i}^{\prime}\boldsymbol{\varepsilon}/T$ is stochastically dominated by a $N\mathopen{}\mathclose{{\left(0,\sigma_{x\varepsilon}^{2}/T}}\right)$, where $\sigma_{x\varepsilon}^{2}=\frac{1-\phi^{2}\phi_{\varepsilon}^{2}}{(1-\phi_{\varepsilon}^{2})(1-\phi\phi_{\varepsilon})^{2}}$. Thus, we have that
$P\mathopen{}\mathclose{{\left(\frac{|\mathbf{x}^{\prime}\boldsymbol{\varepsilon}|}{T}\geq t}}\right)\leq 2e^{-\frac{Tt^{2}}{2\sigma_{x\varepsilon}^{2}}},$
and the union bound yields
[…]
where the second equality follows by setting $t=\sigma_{x\varepsilon}\sqrt{\frac{c_{0}log(n)}{T}}$ for some $c_{0}>2$. Consequently, the inequalities hold with probability at least $1-2e^{-\frac{1}{2}(c_{0}-2)log(n)}$ by setting $\ddot{\lambda}=2\sigma_{x\varepsilon}\sqrt{\frac{c_{0}log(n)}{T}}$, for some $c_{0}>2$. $\square$
A.3 Proof of Theorem 1